1. In the electrolysis of water, how long will it take to produce 1.000 × 102 L of H2 at STP (273 K and 1.00 bar) using an electrolytic cell through which a current of 50.00 mA flows?answer in yr.
1
Expert's answer
2020-05-01T14:23:44-0400

m=E x I x t/26.8.

n= m/M. V=nx22.4.

n=102/22.4= 4.55 moles.

m= n x M = 4.55 x 2= 9.1 g.

t=m x 26.8 / E x I = 9.1 x 26.8/1 x 0.05 = 4 881,42858 hour.

4 881,42858/24 = 203.5 days.